1.     An artificial satellite circles the earth in a circular orbit at a location where the acceleration due to gravity is 9.0ms-2. Determine the orbital period of the satellite.

2.  A satellite of Jupiter has an orbital period of 1.77 days and an orbital radius of . Determine the mass of Jupiter.                                                         

        


1
Expert's answer
2020-12-23T07:32:45-0500
  1. We have artificial satellite 

Let's write a given constants

G = 6.67 * 10-11 ["\frac{m^3*s^2}{kg}"] and "M_{earth}" = 6 * 1024 kg a = 9["\frac{m}{s^2}"]

First, we compute R from a = "\frac{GM}{R^2}", We find R = "\sqrt{\frac{GM}{a}}"

Second we compute v from a = "\frac{v^2}{R}" , which gives v = "\sqrt{aR}" = "\sqrt[4]{GMa}"

Finally, we compute T from this v = "\frac{2\pi R}{T}"

T = "\frac{2\pi R}{v} = 2\pi * \sqrt{GMa} * \frac{1}{\sqrt[4]{GMa}} = 2\pi * \sqrt[4]{\frac{GM}{a^3}}"

= 6.28 * "\sqrt[4]{\frac{6.67 * 10^{-11} * 6 * 10^{24}}{27}}" = 12322[s] = 3.42[hour]

  1. We have T = 1.77days = 1.77 * 86400s and orbital radius R = 7 * 107 m

"T = \frac{2\pi R}{v} = \frac{2\pi R}{\sqrt{aR}} = \frac{2\pi R}{\sqrt{\frac{GM}{R^2}R}} = \frac{2\pi R\sqrt{R}}{\sqrt{GM}} = \frac{6.28 * 7 * 10^7 * \sqrt{7 * 10^7}}{\sqrt{6.67*10^{-11} M}}" = 1.77 * 86400s

"M = 1.8*10^{27}kg"